OpenStudy (anonymous):

of 18 possible books, you plan to take 4 with you on vacation. How many different collections of 4 books can you take?

OpenStudy (anonymous):

you are being asked for the number of ways to choose 4 out of 18 written as \[\binom{18}{4}\] or \[_{18}C_4\] pr even \[^{18}C_4\]

OpenStudy (anonymous):

compute via \[\binom{18}{4}=\frac{\overbrace{18\times 17\times 16\times 15}^{\text{4 term}}}{4\times 3\times 2}\]

OpenStudy (anonymous):

yes, I am using the formula n! / (n-r)!r!

OpenStudy (anonymous):

when I do, I get 1,762,560. that just seems to high

OpenStudy (anonymous):

cancel first, multiply last since the answer is a whole number or just cheat http://www.wolframalpha.com/input/?i=18+choose+4

OpenStudy (perl):

you should get 3060 , after all i said and done

OpenStudy (anonymous):

thank you all

OpenStudy (anonymous):

how about this one 64 numbers draw 6 order doesn't matter, how many combinations?

OpenStudy (anonymous):

There are 56 republicans and 42 democrats and 2 independents how many committees could be formed if each one must contain 3 republicans sand 2democrats?

OpenStudy (anonymous):

help